The reaction 2 NO(g) ---- N2(g)+O2(g) has a Kc value of 2400 at 2000 K. If 0.850 M each of N2 and O2 are initially present in a 3.00-L vessel, calculate the equilibrium of No, N2 and O2


1
Expert's answer
2022-05-15T17:30:03-0400

Ka = products / reactants

= (O2)(N2)/NO

0.850/3 = 0.283 m/L

Initial concentration = 0.283m/L

2400 =2000× 0.283 × X

2400 = 566X

X = 2400/566

= 4.24m/L